A 10 kg block is allowed to slide down a frictionless 600 incline. What is the block's acceleration?

take g = 9.8m/s2

1
Expert's answer
2022-04-27T13:34:55-0400
"a=g\sin\theta\\
=\rm 9.8\: m/s^2\times \sin 60^\circ=8.5\; m/s^2"